PPM2TIFF(1) PPM2TIFF(1) NAME ppm2tiff−createaTIFFfilefromaPPMimagefile SYNOPSIS ppm2tiff[options][input.ppm]output.tif DESCRIPTION ppm2tiffconvertsafileinthePPMimageformattoTIFF.Bydefault,theTIFFimageiscreated withdatasamplespacked(PlanarConfiguration=1),compressedwiththeLempel-Ziv&Welch algorithm(Compression=5),andwitheachstripnomorethan8kilobytes.Thesecharacteristics canbeoverriden,orexplicitlyspecifiedwiththeoptionsdescribedbelow IfthePPMfilecontainsgreyscaledata,thenthePhotometricInterpretationtagissetto1(min-is-black),otherwiseitissetto2(RGB). IfnoPPMfileisspecifiedonthecommandline,ppm2tiffwillreadfromthestandardinput. OPTIONS −cSpecifyacompressionschemetousewhenwritingimagedata:−cnonefornocompression,-cpackbitsforthePackBitscompressionalgorithm,-cjpegforthebaselineJPEG compressionalgorithm,-czipfortheDeflatecompressionalgorithm,and−clzwfor Lempel-Ziv&Welchcompression(thedefault). −rWritedatawithaspecifiednumberofrowsperstrip;bydefaultthenumberofrows/strip isselectedsothateachstripisapproximately8kilobytes. −RMarktheresultantimagetohavethespecifiedXandYresolution(indots/inch). SEEALSO tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) October15,1995 1